OpenStudy (anonymous):

Please help! I only have about 20 minutes for this problem! A circle has its center at (2, -4) and a radius of 6 units. What is the equation of the circle?

OpenStudy (♪chibiterasu):

Do you know how to form an equation for a circle?

OpenStudy (anonymous):

Not really, I didn't understand that part in my math lesson quite well. It's the only thing I'm struggling with.

OpenStudy (mertsj):

Why do you only have 20 minutes for this question?

OpenStudy (anonymous):

I'm going somewhere in 20 minutes but this was a last minute thing. :/

OpenStudy (mertsj):

What is the (h,k) form of the equation of a circle?

OpenStudy (♪chibiterasu):

The circle's center is not at the origin so we will be using the h, k equation. \[(x - h)^2 + (y - k)^2 = r^2\] You just . . . fill it in. This is really all you need to know. If the center is at the origin just use: \[x^2 + y^2 = r^2\]

OpenStudy (♪chibiterasu):

You fill in for h and k, not for x and y.

OpenStudy (♪chibiterasu):

Remember to alternate the signs when necessary!

OpenStudy (anonymous):

(x - h)^2 + (y - k)^2 = r^2 (h, k) = center and r = radius so for when its center is at (2, -4) and it has a radius of 6 plug in h=2 , k=-4 , and r=6 into the circle equation ?? That correct?

OpenStudy (♪chibiterasu):

yeah

OpenStudy (♪chibiterasu):

Remember that inputting negatives will switch the - to a +